OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

OpenStudy (anonymous):

3x-8=7 (combine same terms) +8 +8 (Add 8 both sides) 3x=15 (Now just divide!) __ __ 3 3 X=5! (Vwalla! ANSWER! MEDAL! :D )